# LLM-SQL-Solver: Can LLMs Determine SQL Equivalence?

### v3

@@ 1 Introduction
Structured Query Language (SQL) it is the standard programming language to query data systems and it is one of the most popular programming language Stackscale . Since SQL is not a procedural language, one may write different SQL queries that are semantically equivalent. As a result, determining the equivalence of SQL queries is a fundamental task in many practical important applications ranging from query optimizers in database systems Pirahesh et al. (1992); Bai et al. (2023); Zhou et al. (2021); Samwel et al. (2018) to evaluating the translation from user text-to-SQL Zelle and Mooney (1996); Zhong et al. (2020b); Khatry et al. (2023); Floratou et al. (2024a). However, it has been shown by the data management research community that checking the equivalence of two SQL queries is in general undecidable Abiteboul et al. (1995). While there are theoretical works to check equivalence under different assumptions Chandra and Merlin (1977); Sagiv and Yannakakis (1980); Cohen et al. (1999); Chu et al. (2018) and implemented prototypes such as Cosette Chu et al. (2017), they only support limited SQL operations and such projects are no longer maintained due to lack of experts. For instance, Cosette does not support commonly found operations like EXISTS, IN, queries with constant expressions, etc.; and Cosette is no longer maintained due to the lack of staff. Query optimizers and query processing engines in data systems still primarily rely on heuristic query rewriting rules Begoli et al. (2018). Exploring and writing these rules correctly require strong domain expertise and need many engineering hours.
@@
Moreover, due to the difficulty in evaluating the equivalence of two SQL queries, Text-to-SQL literature Zhong et al. (2017); Yu et al. (2018); Wang et al. (2019); Lee et al. (2021); Li et al. (2023) primarily use exact match (exact SQL string match) or execution match (compare the execution outputs) as the evaluation metrics. However, exact string match may overlook many equivalent SQL queries (there are many ways different ways to write the same SQL query), and execution match may misclassify different SQL queries as equivalent due to the lack of counterexamples in the underlining database instance. We note that there have been a few attempts that explore using semantic equivalence as the evaluation metrics Zhong et al. (2020a). These methods assess query semantic equivalence by generating numerous fuzzy database instances, leading to significant computational and storage demands. Specifically, evaluating a thousand pairs of SQL queries requires at least 75 minutes and consumes about 3 GB of storage space for creating these fuzz database instances Zhong et al. (2020a).
@@
Recent advancements in large language models (LLMs) have shown great capability in following instructions and in generating decisions that align with human judgement Bubeck et al. (2023); Chiang and Lee (2023); Yan et al. (2023a); Zhang et al. (2023b). For instance, It has been shown that GPT models are capable of performing first order logic reasoning Saparov and He (2022), detecting implicit contents Huang et al. (2023); Zhang et al. (2023c), and solving complex logical problems Feng et al. (2023); Yao et al. (2023). In this paper, we investigate whether LLM can act as a judge to determine the equivalence of SQL queries. Hence, in this paper, we would like to mainly answer the following questions:
@@ 2.1 Motivations
Hence, real-world data systems need tools to avoid such pitfalls. While proving the semantic equivalency between two SQL queries has been proved to be undecidable Trahtenbrot (1963); Libkin (2004); Sagiv and Yannakakis (1980), we believe LLMs may offer new opportunities to help data engineers in scrutinizing their SQL optimization logic and potentially identify flaws accordingly. For instance, when we present the COUNT BUG to ChatGPT, ChatGPT is able to correctly identify that these two SQL queries are not equivalent.
@@ 2.1.1 Two Notions of SQL Equivalence
Definition 2.1 (Semantic Equivalence).
@@
Definition 2.2 (Relaxed Equivalence).
Two SQL queries Q1 and Q2 are said to be relaxed equivalent if one may edit a trivial portion of the SQL queries while preserving the same logic to make Q1 and Q2 semantically equivalent.
2.2 Why LLM-SQL-Solver
First, LLM-SQL-Solver (semantic notion) can be directly used to find system bugs (e.g., COUNT BUG) and help engineers validate their SQL rewriting code bases. Second, since query optimizer relies on heuristic rules to find equivalent SQL forms, it may not explore all rewriting possibilities and hence, the rewritten query may lead to sub optimal performance. We believe that, with LLM-SQL-Solver, query optimizers can explore the vast universe of equivalent rewritten queries and then find the query plan with near optimal performance.
For text-to-SQL tasks, we believe that when comparing the gold SQL query (the correct answer) with the generated SQL query, the current most popular metric, execution match, does not reflects the human perspective and fails to capture a user’s opinion on the generated SQL. For instance, a common reasoning for failures in execution match is that the generated SQL selects fewer/extra columns than the gold SQL. In fact, a recent work has explicitly added "do not select extra columns" in its prompt to improve the execution accuracy Dong et al. (2023). However, from a user’s perspective providing extra columns may not reduce the quality of the answer. Also, there are many other similar reasons that can lead to false execution matches but not necessarily indicate a lower quality of the generated SQL query (such as the column order in selection, case sensitivity, duplicate entries, and columns in different tables share the similar contents). Moreover, execution matching fails to correctly identify logically different queries when the underlying database does not contain any counter examples, even when hundreds of human hours were consumed to create the benchmark Yu et al. (2018). Lastly, a fundamental limitation of execution accuracy is its need to query the underlying database. This is particularly pertinent as the text-to-SQL service provider usually does not have ownership of the user’s database and the user’s database may also contain sensitive information. As a result, it would be desirable to have a SQL solver determining the relax notion of equivalence without constructing sophisticated databases and without accessing any stored data. Hence, in additional to investigating the capability of foundational models in identifying semantic equivalence, we further explore whether LLM-SQL-Solver (relaxed notion) can be used as a better metric for text-to-SQL benchmarks.
@@ 3 Proposed Method
Motivated by recent methods Deng et al. (2023); Yao et al. (2023); Wei et al. (2022), we present our prompting techniques to obtain high quality judgments from Open-AI GPT models. Our prompting technique adheres to key aspects widely adopted in NL2SQL works Li et al. (2023); Khatry et al. (2023); Liu et al. (2023); Taori et al. (2023b): 1) Clear Layout: We provide the instruction at the beginning, followed by the two SQL queries that need to be compared, and end by providing the database schema. 2) Clear Context: We provide the database schema and explicitly included the primary and foreign key constraints, since they are crucial context information for deciding SQL equivalence.
For semantic equivalence, we also added a short definition of SQL semantic equivalence (i.e., Note: Two SQL queries are semantically equivalent if and only if they yield identical results for all possible databases.). In addition, we first ask GPT to generate a counter example if there exists one and then decide if two SQL queries are equivalent. We call this prompt the Counter Example prompt. However, as shown in Figure 1, the counter example prompt leads to sub-optimal results, especially when GPT models only scratched the surface and observed the obvious (similarity between two SQL queries in structure and logic). This finding aligns with the prior works Wang et al. (2023); Zheng et al. (2023) which demonstrated that GPT models exhibit bias toward ignoring small differences when determining SQL equivalence and exhibit limited reasoning capability.
@@
To overcome this observed phenomenon, we introduce our proposed prompting technique in detail to obtain better responses from large language models. In complex problem solving, people always start with a simple example and then introduce more complex elements to understand the full picture. Motivated by this simple intuition, we incorporate this idea and proposed the Miniature and Mull prompting technique. In particular, we first ask GPT to execute both queries on a simple database and then see if it can modify the database such that the two queries output different results (i.e., a counter example). As shown later in the experimental section, this approach can directly enhance the LLM’s response accuracy in determining SQL semantic equivalence. We believe that this prompting technique has wide applications in comparing the equivalence of two codes, plans, and logic.
For relaxed equivalence, we designed the Explain and Compare prompt by asking the LLMs to perform multi-step reasoning before determining if the two SQL queries are equivalent. This prompt follows the main ideas from the chain-of-thoughts Wei et al. (2022). We first prompt the LLMs to explain two SQL queries and then ask it to determine if these two queries contain significant difference in logic, as shown in Figure 1. We observe that LLMs are inclined to label two SQL queries as equivalent when they appear similar in structure. As a result, it is crucial for LLMs to be able to first spot the differences between these two queries and then apply this knowledge in considering if these two SQL queries are relaxed equivalent.
@@ 4 Evaluations
We evaluate the capabilities of popular large language models in identifying semantic and relaxed equivalence between a pair of SQL queries. In particular, we showcase the experimental results using the popular GPT 3.5 turbo (gpt-3.5-turbo) and GPT 4 turbo (gpt-4-1106-preview) models. The models are accessed via OpenAI API.
Hyper-parameter: We use zero demonstrations in our prompts. In addition, in the API call, we have set temperature to 0 which implies greedy sampling is used to generate the output; The max tokens are set to 3000; and the stopping tokens are set to "Answer:" for all experiments.
@@ 4.1 Data Source
First, we compare our proposed Miniature&Mull prompt with the CounterExample prompt on i) Non-equivalent Queries Chu et al. (2017) and ii) Equivalent Queries Zhong . These two data sources include pairs of SQL queries that have been expertly labeled as semantically non-equivalent or equivalent.
Nonequivalent Queries Chu et al. (2017): Four challenging and one easy questions, including the COUNT BUG, one exam question from CSE344 at the University of Washington, two bugs users found on SQL solver, and a simple hello world example.
Equivalent Queries Zhong et al. (2020a): A hundred pairs of expert labeled equivalent queries.
Next, we showcase results on the alignment of our proposed LLM-SQL-Solver judgements with human preferences on semantic and relaxed notions using the Spider benchmark Yu et al. (2018). The Spider benchmark contains 1034 development questions covering 20 databases in different domains. The Spider SQL queries are classified into four levels of complexity: easy, medium, hard, and extrahard (extra). We evaluate the judgments of LLM-SQL-Solver against execution accuracy (EX) based on human preferences. In all experiments, we use the open sourced EX evaluation code from GitHub 111https://github.com/THU-BPM/ChatGPT-sql.

### v4

@@ 1 Introduction
Structured Query Language (SQL) it is the standard programming language to query data systems and it is one of the most popular programming language Stackscale . Since SQL is not a procedural language, one may write different SQL queries that are semantically equivalent. As a result, determining the equivalence of SQL queries is a fundamental task in many practical important applications ranging from query optimizers in database systems Pirahesh et al. (1992); Bai et al. (2023); Zhou et al. (2021); Samwel et al. (2018) to evaluating the translation from user text-to-SQL Zelle and Mooney (1996); Zhong et al. (2020b); Khatry et al. (2023); Floratou et al. (2024a). However, it has been shown by the data management research community that checking the equivalence of two SQL queries is in general undecidable Abiteboul et al. (1995). While there are theoretical works to check equivalence under different assumptions Chandra and Merlin (1977); Sagiv and Yannakakis (1980); Cohen et al. (1999); Chu et al. (2018); Wang et al. (2022b) and implemented prototypes such as Cosette Chu et al. (2017), they only support limited SQL operations and such projects are no longer maintained due to lack of experts. For instance, Cosette does not support commonly found operations like EXISTS, IN, queries with constant expressions, etc.; and Cosette is no longer maintained due to the lack of staff. Query optimizers and query processing engines in data systems still primarily rely on heuristic query rewriting rules Begoli et al. (2018). Exploring and writing these rules correctly require strong domain expertise and need many engineering hours.
@@
Moreover, due to the difficulty in evaluating the equivalence of two SQL queries, Text-to-SQL literature Zhong et al. (2017); Yu et al. (2018); Wang et al. (2019); Lee et al. (2021); Li et al. (2023) primarily use exact match (exact SQL string match) or execution match (compare the execution outputs) as the evaluation metrics. However, exact string match may overlook many equivalent SQL queries (there are many ways different ways to write the same SQL query), and execution match may misclassify different SQL queries as equivalent due to the lack of counterexamples in the underlining database instance. We note that there have been a few attempts that explore using semantic equivalence as the evaluation metrics Zhong et al. (2020a). SQL queries in text-to-SQL benchmarks often include constructs such as EXISTS, UNION, UNION ALL, and database schema doesn’t no exclude NULL values, making off-the-shelf SQL provers (e.g., Cosette Chu et al. (2017) and WeTune Wang et al. (2022b)) unsuitable for direct use. Consequently, researchers have turned to fuzzy testing, which assesses query semantic equivalence by generating numerous fuzzy database instances. However, this approach imposes substantial computational and storage demands. Specifically, evaluating a thousand pairs of SQL queries requires at least 75 minutes and consumes about 3 GB of storage space for creating these fuzz database instances Zhong et al. (2020a). This substantial data requirement arises from a common issue in execution-based approaches: incorrectly classifying semantically inequivalent queries as equivalent due to the not having a counter-example included. For example, when the underlying database lacks tuples matching the predicate and join conditions, two semantically inequivalent queries can both produce an empty tuple as output. As a result, these queries will appear to match, regardless of their actual differences.
@@
Recent advancements in large language models (LLMs) have shown great capability in following instructions and in generating decisions that align with human judgment Bubeck et al. (2023); Chiang and Lee (2023); Yan et al. (2023); Zhang et al. (2023b). For instance, It has been shown that GPT models are capable of performing first order logic reasoning Saparov and He (2022), detecting implicit contents Huang et al. (2023); Zhang et al. (2023c), and solving complex logical problems Feng et al. (2023); Yao et al. (2023). In this paper, we investigate whether LLM can act as a judge to determine the equivalence of SQL queries. Hence, in this paper, we would like to mainly answer the following questions:
@@
The paper is organized as follows: Section 2 motivates the need of LLMs in determining SQL equivalence and discusses the background information of two notions of SQL equivalence and LLMs basics. Section 3 introduces two prompting methods, Miniature and Mull and Explain and Compare, to improve LLMs’ ability to assess SQL queries semantic and pragmatic equivalence. Section 4 presents experimental results on assessing SQL equivalence using GPT-3.5-Turbo and GPT-4o respectively, applying our two proposed prompting techniques on three datasets from the Spider benchmark. Section 5 discusses the implications of the proposed LLM-SQL-Solver for SQL query validation and text-to-SQL evaluation, while also highlighting potential areas of improvement for our approach. Finally, Section 6 summarizes our contributions and concludes this work.
@@ 2.1 Motivations
Hence, real-world data systems need tools to avoid such pitfalls. While proving the semantic equivalency between two SQL queries has been proved to be undecidable Trahtenbrot (1963); Libkin (2004); Sagiv and Yannakakis (1980), we believe LLMs may offer new opportunities to help data engineers in scrutinizing their SQL optimization logic and potentially identify flaws accordingly. For example, when we presented the COUNT BUG to ChatGPT, it was able to correctly identify that the two given SQL queries were not equivalent, effectively pointing out the underlying issue. Highlighting the potential of LLMs to complement traditional approaches by offering new insights.
@@
Moreover, in text-to-SQL generation, many natural language questions are inherently ambiguous Katsogiannis-Meimarakis and Koutrika (2023); Floratou et al. (2024b), allowing for multiple valid interpretations and corresponding answers Min et al. (2020). From a user’s perspective, different SQL queries that yield the same logical answer to a question are often considered equivalent. The primary objective of text-to-SQL systems is to align with human preferences and deliver outputs that meet user expectations. However, the commonly used execution match metric, which assumes a single correct SQL query, fails to account for these ambiguities. For example, consider a ground truth SQL query like "SELECT shipdate FROM Supply" and a predicted SQL query such as "SELECT shipdate, COUNT(shipdate) FROM Supply". Although these queries produce different output tables (making them execution unmatched), they convey the same underlying information. In such cases, a user might find the predicted query acceptable, even though it fails the strict execution match criterion. The limitations of execution match highlights the need for alternative approaches to evaluate the quality of text-to-SQL generation. On the other hand, large language models (LLMs) are often trained using Reinforcement Learning with Human Feedback (RLHF), which aligns their outputs more closely with human preferences and judgments Bai et al. (2022). We are particularly interested in exploring whether LLMs can provide a more nuanced assessment by identifying and validating diverse SQL queries that share the same intent, thereby achieving better alignment with human preferences.
@@ 2.1.1 Two Notions of SQL Equivalence
Definition 2.1 (Semantic Equivalence Codd (1970)).
@@
Definition 2.2 (Pragmatic Equivalence Baker (2018)).
Two SQL queries Q1 and Q2 are said to be pragmatic equivalent if they produce outputs that convey the same practical or meaningful information for the intended user, even if the outputs differ in format, structure, or additional details.
Semantic equivalence is a well understood concept in database community. In contrast to semantic equivalence, pragmatic equivalenceBaker (2018) is originated in natural language translation communities, focuses on maintaining the intended meaning or effect of an expression rather than its literal structure. For example, in natural language translation, a direct translation of the English idiom ’It rains cats and dogs’ into another language would fail to convey the intended meaning, as the phrase is nonsensical. Instead, the pragmatically equivalent translation would be ’rain heavily,’ which conveys the same idea of intense rainfall in a way that is meaningful to a non-English-speaking audience, without referencing cats or dogs. Similarly, in the context of SQL, pragmatic equivalence considers whether two queries achieve the same user intent or provide comparable practical information, even if their outputs or structures differ. This concept is particularly relevant in text-to-SQL systems, where multiple distinct yet valid queries can effectively fulfill the same user intent.
Based on the semantic definition of SQL equivalence (Definition 2.1), the Count bug is indeed a flaw for query engines and should be remedied immediately. On the other hand, we argue that based on Definition 2.2, which reflects the logical perspective and aligned with human perception, the two SQL queries in the Count Bug example are pragmatic equivalent. The fact that even experienced experts failed to identify this subtle difference indicates humans often recognize equivalence based on the more relaxed pragmatic notion. This observation motivates us to explore whether LLMs can serve as a more effective metric for evaluating text-to-SQL tasks, aligning better with human judgment.
2.2 Large Language Models Basics
Large language models (LLMs) are trained on extensive, web-scale datasets and leverage attention mechanisms Vaswani et al. (2017) combined with Reinforcement Learning with Human Feedback (RLHF) to generate high-quality, contextually relevant responses to input prompts. These models have demonstrated exceptional capabilities in solving complex tasks across diverse domains, such as natural language-to-SQL translation Pourreza and Rafiei (2023); Gao et al. (2023), data system optimizations Zhou et al. (2023), and data discovery Trummer (2023); Fernandez et al. (2023).
A defining characteristic of LLMs is their ability to function as few-shot learners Brown et al. (2020). One widely used approach to enhance their accuracy is in-context learning (ICL). Through ICL, LLMs utilize example demonstrations provided within the input prompt to improve task performance dynamically. Morevoer, when prompted with specific instructions and requirements, such as generating output in JSON format, LLMs adapt their responses to align with these constraints, ensuring accurate and tailored results.
@@ 3 Proposed Method
Motivated by recent methods Deng et al. (2023); Yao et al. (2023); Wei et al. (2022), we present our prompting techniques to obtain high quality judgments from Open-AI GPT models. Our prompting technique adheres to key aspects widely adopted in text-to-SQL works Li et al. (2023); Khatry et al. (2023); Liu et al. (2023); Taori et al. (2023): Clear Layout: We provide the two SQL queries that need to be compared at the beginning, then followed with the database schema, and end with the task (e.g., semantic equivalence or pragmatic equivalence) and instructions. Clear Context: We provide the database schema and explicitly included the primary and foreign key constraints, since they are crucial context information for deciding SQL equivalence. Prompt style: We adopted the code-representative approach, which has proven highly effective for database-related tasks Gao et al. (2023). An example prompt is illustrated below:
For evaluating semantic equivalence, we propose two distinct prompts. One approach utilizes a chain-of-thought (CoT) prompt, where the task and instructions explicitly guide the model to determine semantic equivalence by reasoning step-by-step. While the chain-of-thought prompt is generally effective, as shown in Figure 1, it often yields suboptimal results. This is particularly evident when language models merely scratch the surface, focusing on obvious similarities in the structure and logic of SQL queries rather than nuanced differences. This observation aligns with prior studies Wang et al. (2023); Zheng et al. (2023), which highlight that language models tend to overlook subtle distinctions when assessing equivalence and demonstrate limited reasoning capabilities in such contexts.
@@
To overcome this observed phenomenon, we introduce our proposed prompting technique in detail to obtain better responses from large language models. In complex problem solving, people always start with a simple example and then introduce more complex elements to understand the full picture. Motivated by this simple intuition, we incorporate this idea and proposed the Miniature and Mull prompting technique. In particular, we first ask LLMs to execute both queries on a simple database. If the outputs of the two SQL queries are identical, we then instruct LLMs to adjust the simple database (e.g., updating rows) and observe how the adjustments affect the outputs. If the modified database causes the two queries to produce different results, it indicates semantic inequivalence (i.e., a counterexample exists).
As shown later in the experimental section, this approach can directly enhance the LLM’s response accuracy in determining SQL semantic equivalence. We believe that this prompting technique has wide applications in comparing the equivalence of two codes, plans, and logic.
In addition, we have added a specific instruction in both CoT and Miniature & Mull prompt: "The SQL predicate is case-sensitive when comparing string values". We have observed that LLMs often assume databases are case-insensitive. While many databases, such as SQL Server Mistry and Misner (2014) and MySQL DuBois (2013), are indeed case-insensitive by default, the Spider benchmark employs SQLite, which is case-sensitive by default. To address this, both the CoT and Miniature & Mull prompts explicitly specify that the database is case-sensitive when comparing string values.
For assessing pragmatic equivalence, we observed that LLMs often exhibit a tendency to label two SQL queries as equivalent based primarily on their structural and semantic similarity. This bias can lead to erroneous conclusions, as structural resemblance does not necessarily imply logical equivalence. To address this challenge, it is essential for LLMs to first identify and articulate the differences between the queries and then leverage this understanding to evaluate whether the queries are pragmatically equivalent. As a result, we designed the Explain and Compare prompt by asking the LLMs to perform multi-step reasoning before determining if the two SQL queries are equivalent. This prompt follows the main ideas from the chain-of-thoughts Wei et al. (2022). We first prompt the LLMs to explain two SQL queries to understand their high-level objectives and then ask LLMs to determine if these two queries contain aligned logic or significant differences, as shown in Figure 1. This structured approach mitigates the LLMs’ inclination to focus solely on surface-level similarities. By requiring them to articulate reasoning steps, the prompt ensures a deeper evaluation of equivalence.
Moreover, we added specific instruction in Explain & Compare prompt regarding NULL values and case-sensitivity: "Database contains no NULL values". We have observed that LLMs often consider COUNT(*), COUNT(primary key), and COUNT(DISTINCT primary key) to be different. The reason is that LLMs consider cases when NULL values can be inserted into a column with primary key constraint. In fact, the SQLite databases, by default, allow null values to be inserted to primary key column. However, in many database primary key constraints explicitly states that null values are not allowed Oracle . Although LLMs correctly consider that having NULL values in the primary key column may lead to different results (note that COUNT(*) considers null values as unique entries, where as COUNT(primary key) ignores rows with null values), these are not significant issue and does not affect the meaning of the query.
@@ 4 Evaluations
We evaluate the capabilities of popular large language models in identifying semantic and pragmatic equivalence between a pair of SQL queries. In particular, we aim to answer the following questions:
Is the Miniature & Mull prompt effective in accurately determining semantic equivalence?
Does LLMs using the Explain & Compare prompt achieve better alignment with human preferences than execution match?
We showcase the experimental results using the popular GPT-3.5-Turbo (gpt-3.5-turbo) and GPT-4o (gpt-4o) models. The models are accessed via OpenAI API.
Hyper-parameter: In the API call for all experiments, we have set temperature to 0 which implies greedy sampling is used to generate the output; The max tokens are set to 4096.
Few shot demonstrations. In addition to the zero-shot setting, we also investigate the in-context learning capabilities of LLMs in determining SQL equivalence. This is achieved by providing five static few-shot examples as part of the input context to guide the model’s reasoning process. For the semantic equivalence, the five static examples include three semantically equivalent SQL queries and two semantically inequivalent SQL queries. Similarly, for pragmatic equivalence, the five static examples consist of three equivalent and two inequivalent query pairs. The equivalent examples feature queries that share the same intent despite differences in their structure. For instance, queries like "SELECT birth_country FROM player" and "SELECT birth_city FROM player" are considered equivalent as both aim to determine the birth location, albeit at different levels of granularity.
@@ 4.1 Data Source
First, we compare our proposed Miniature & Mull prompt with the CoT prompt on i) Calcite test suites Begoli et al. (2018); Ding et al. (2023) and ii) Inequivalent Queries Gao et al. (2023). These two data sources include pairs of SQL queries that have been expertly labeled as semantically equivalent or non-equivalent.
Next, we present the results evaluating the alignment between our proposed LLM-SQL-Solver’s pragmatic equivalence judgments and human preferences. To conduct this assessment, we selected 70 pairs of SQL queries, comprising gold-standard SQL queries from the Spider benchmark Yu et al. (2018) and predicted SQL queries generated by LLMs Gao et al. (2023); Pourreza and Rafiei (2023); Liu et al. (2023). Each SQL pair was independently reviewed and labeled by at least three human experts proficient in SQL. The final label for each pair was determined through a majority voting process, ensuring consensus among the reviewers. This approach provides a robust benchmark for comparing the LLM’s pragmatic equivalence evaluations with human judgment. In addition to the human preference, we also provide the execution accuracy (EX) by executing these 70 queries over Spider databases. In all experiments, we use the open sourced EX evaluation code from GitHub ***https://github.com/THU-BPM/ChatGPT-sql.
Calcite test suites: It contains 232 pairs of equivalent queries (verified by expert) in the Calcite test suites.
Semantically Inequivalent Queries: It contains 180 pairs of semantic inequivalent queries. One of the query is the gold SQL in spider benchmark Yu et al. (2018) and the other is a predicted SQL query from DAIL-SQL Gao et al. (2023). Because they have different execution outputs, the databases in spider benchmark serve as counter examples.
Pragmatic Equivalence Queries: It contains 70 pairs of SQL queries with human preference label. Each pair has been independently reviewed and labeled by at least three human experts to capture human preference judgments.